MONT 105N – Analyzing Environmental Data February 15, 2019 Rhopography Study Guide

Background

Our reading for class on Friday, February 22 is a section of the chapter titled “Rhopography” from the book Looking at the Overlooked: Four Essays on Still Life Painting by Norman Bryson. This is available as a scanned .pdf on the course homepage. This is a challenging text that will require some careful attention and thought. To help you to get the most of this (and our museum visit the following Sunday, February 24), here are some study questions.

Questions

1. Who is the author and what is his background?

2. What does the author mean by the terms “rhopography” and “megalography”

3. The title of Bryson’s book is Looking at the Overlooked: Four Essays on Still Life Painting. What do you think this title means?

4. Bryson positions still life painting in relation to other painting genres such as portraiture and history painting. How does he make this comparison?

5. Bryson uses the paintings of Juan Sánchez Cotán as “examples of rhopography’s potential for overturning the scale of human importance?” Discuss how.

6. What are bodegónes?

7. According to Bryson, Cotán’s work engages “the eye” in a powerful and particular way. Discuss how.

8. How are Cotán’s paintings “hyperrealistic” rather than merely “realistic”? What is the difference?

9. Describe how Bryson draws a comparison between the Spiritual Exercises of St. Ignatius and Cotán’s paintings.

10. How are Cotán’s paintings “anorexic” according to Bryson?

11. Describe how Bryson describes Cotán’s paintings in mathematical terms. Note: The footnote 6 on page 69 was not copied. It says: “To be more precise about Cotán’s calculus: if the estimated centres of gravity for each fruit/vegetable are joined, the result is what is known as an exponential decay curve.” Is that different from a “logarithmic curve?” How?

12. According to Bryson, what is the ultimate aim of Cotán’s mathematical approach to composition and his painting technique? Do you agree with this argument?